# Reward Model Overoptimisation in Iterated RLHF

### v1

@@ Abstract
Reinforcement learning from human feedback (RLHF) is a widely used method for aligning large language models with human preferences. However, RLHF often suffers from reward model overoptimisation, in which models overfit to the reward function, resulting in non-generalisable policies that exploit the idiosyncrasies and peculiarities of the reward function. A common mitigation is iterated RLHF, in which reward models are repeatedly retrained with updated human feedback and policies are re-optimised. Despite its increasing adoption, the dynamics of overoptimisation in this setting remain poorly understood. In this work, we present the first comprehensive study of overoptimisation in iterated RLHF. We systematically analyse key design choices - how reward model training data is transferred across iterations, which reward function is used for optimisation, and how policies are initialised. Using the controlled AlpacaFarm benchmark, we observe that overoptimisation tends to decrease over successive iterations, as reward models increasingly approximate ground-truth preferences. However, performance gains diminish over time, and while reinitialising from the base policy is robust, it limits optimisation flexibility. Other initialisation strategies often fail to recover from early overoptimisation. These findings offer actionable insights for building more stable and generalisable RLHF pipelines.
@@ 1 Introduction
Reinforcement learning from human feedback (RLHF) has become the standard method for aligning large language models with human preferences (ziegler2020finetuninglanguagemodelshuman, ; ouyang2022traininglanguagemodelsfollow, ; bai2022traininghelpfulharmlessassistant, ). However, RLHF faces a critical vulnerability: reward model overoptimisation (gao2022scalinglawsrewardmodel, ). As fine-tuning progresses, models learn to overfit to the trained reward function - achieving high scores without genuinely satisfying human intent. This creates brittle policies that exploit loopholes rather than developing robust behaviours, leading to systems that appear aligned during training but fail catastrophically when deployed. Iterated RLHF represents a promising approach to combat this problem. By repeatedly collecting new preferences on the latest policy outputs, retraining the reward model, and fine-tuning the policy (bai2022traininghelpfulharmlessassistant, ; xiong2024iterative, ), practitioners aim to iteratively close the gap between proxy and true reward. Despite its widespread adoption in industry (ziegler2020finetuninglanguagemodelshuman, ; ouyang2022traininglanguagemodelsfollow, ; bai2022traininghelpfulharmlessassistant, ), it remains uncertain whether iterated RLHF genuinely resolves overoptimisation, merely postpones the inevitable exploitation of the reward model akin to persistent adversarial policies (gleave2021adversarial, ), or perpetuates a recurring cycle of overoptimisation in different forms (singhal2024longwaygoinvestigating, ).
@@
Figure 1: The Iterated RLHF framework performs multiple rounds of preference data collection, reward modelling, and policy optimisation. Our research reveals three design choices that dramatically impact performance: (1) how preference data is managed across iterations, (2) which reward function formulation to optimise, and (3) how policies are initialised at each stage. Effectively configuring these elements can significantly reduce overoptimisation while improving true reward.
Figure 2: Our design choices for preference data management, reward function formulation, and policy initialisation significantly impact performance across iterations. Concatenating preference data yields dramatic improvements in true reward scores, while equal-sampling with last-policy initialization performs worst (see Section 5). These findings offer concrete guidance for practitioners implementing robust RLHF systems.
@@
  • We provide guidelines for practitioners implementing iterated RLHF, including specific recommendations for preference data management, reward function selection, and policy initialisation strategies.
Using a gold-standard reward model to simulate human labellers (coste2024reward, ; gao2022scalinglawsrewardmodel, ) on the AlpacaFarm dataset (alpaca, ) and working exclusively with open-source models, our experiments yield several key insights: Reward models become increasingly robust across iterations, leading to higher gold reward scores (Figure 2). Performance gains diminish after three iterations for most methods. Concatenating preference data across iterations dramatically outperforms other approaches. Small but persistent overoptimisation remains after four iterations regardless of design choices.
Our results demonstrate that while iterated RLHF significantly improves reward model robustness, it does not fully eliminate overoptimisation. This underscores the need for continued research into more robust alignment methods that can withstand sophisticated specification gaming (krakovna2020specification, ) by increasingly capable language models.
@@ 2 Related work
RLHF is the standard for aligning large language models to human preference data. The iterated approach has been first used by the authors in bai2022traininghelpfulharmlessassistant to fix robustness and calibration issues, attributed to lack of data in the high score regime and has since gained in popularity ramé2024warpbenefitsweightaveraged ; xiong2024iterative ; Ye2024OnlineIR ; adolphs2022cringe ; Dong2024RLHFWF ; yuan2024selfrewarding . Besides training on newly collected preferences, an iterated scheme to train reward models from synthetically generated preference data has been proposed in wang2024selftaughtevaluators and shown to improve performance on the reward model benchmark RewardBench lambert2024rewardbenchevaluatingrewardmodels , but the authors focus on iterated training of an evaluator and do not study overoptimisation nor the design choices we consider. In the context of Direct Preference Optimisation (DPO) rafailovDirectPreferenceOptimization2023 offline, online and hybrid approaches repeatedly collecting new preference data have been investigated mostly in terms of sample efficiency (xiong2024iterative, ; das2024provably, ; muldrew2024active, ; mehtaSampleEfficientReinforcement2023, ). More broadly iterated methods have been investigated for machine teaching (liu2017iterative, ) and to resolve feedback loops caused by model deployment in supervised learning (pmlr-v119-perdomo20a, ) and also performative RL (Mandal2022PerformativeLearning, ).
Overoptimisation is a common issue in RL, and evidence of this has been frequently reported in the RLHF literature (ziegler2020finetuninglanguagemodelshuman, ; stiennon2022learningsummarizehumanfeedback, ; gao2022scalinglawsrewardmodel, ; singhal2024longwaygoinvestigating, ). A promising method to mitigate overoptimisation is using reward model ensembles combined with conservative optimisation as proposed by coste2024reward . Several works further explore reward model ensembles in RLHF eisenstein2023helping ; lou2024uncertainty . Notably, in ramé2024warm the authors introduce weight averaged reward models (WARM) alleviating the inference cost of multiple reward models during training. Following coste2024reward and gao2022scalinglawsrewardmodel in tackling reward model overoptimisation, several works propose alternative approaches including reward model distillation fisch2024robust , hidden state regularisation Yang2024RegularizingHS , and more Yang2024BayesianRM ; Miao2024InfoRMMR ; Liu2024ProvablyMO ; gorbatovski2024learnreferencemodelreal . One commonly reported mode of overoptimisation is length bias (singhal2024longwaygoinvestigating, ; Park2024DisentanglingLF, ), which can be tackled by disentangling reward signals related to response length from content quality (chen2024odin, ). To the best of our knowledge, the literature lacks a systematic investigation into overoptimisation in iterated Reinforcement Learning from Human Feedback (RLHF). Such an investigation is not only necessary but also fundamentally important for a deeper understanding and meaningful improvement of fine-tuning methods based on this technique.
@@ 3 Iterated Reinforcement Learning from Human Feedback
In this section, we first outline the process of a single iteration of RLHF and then extend it to the iterated framework. The RLH pipeline consists of the following three steps: 1. Collection of a preference dataset; 2. Reward model training; 3. Policy optimisation on the reward model. Though not an integral part of the RLHF pipeline, it is common in practice for step 1 to be preceded by supervised fine-tuning on labelled examples.
@@ 3.1 Single-iteration RLHF
Preference data collection. We begin with a policy $\pi^{sft}$, obtained by supervised fine-tuning (in practical terms, a policy checkpoint), and use it for collecting preference data. More specifically, the preference dataset $\mathcal{D}$ consists of tuples $\{x_{i},y_{i,0},y_{i,1},p_{i}\}$ for $i=1,...,N$, where $x_{i}\in\mathcal{X}$ is a prompt or instruction, $y_{i,j}\sim\pi^{sft}(\cdot|x_{i})$ with $j=0,1$ are two synthesised responses from the supervised fine-tuned policy checkpoint $\pi^{sft}$, and $p_{i}$ is the preference label indicating whether $y_{i,0}$ is preferred over $y_{i,1}$. Similarly to coste2024reward ; gao2022scalinglawsrewardmodel , the preference labels $p_{i}$ are simulated with a gold reward model denoted by $R^{\star}$, which is significantly larger in terms of parameter size than the proxy reward models, as an approximation for human labellers that would be used in the real RLHF process instead.
Reward model training. The proxy reward model $R_{\phi}$ is initialised from model checkpoint $R^{\text{init}}$, with a randomly initialised prediction head, and subsequently trained by minimizing the cross-entropy loss on the preference dataset $\mathcal{D}$. It is standard to use the Bradley-Terry model bradley1952rank , under which the probability of preferring the answer $y_{0}$ over $y_{1}$ given prompt $x$ is given by
@@
Policy optimisation. Having trained the proxy reward model $R_{\phi}$, the policy $\pi_{\theta}$ is initialised from $\pi^{sft}$ and then fine-tuned to optimise $R_{\phi}$. This is commonly achieved with the proximal policy optimization (PPO) algorithm schulman2017proximal . In order to prevent overoptimisation of the proxy reward model and regularise $\pi_{\theta}$ to not diverge too drastically from its initialisation, a Kullback-Leibler divergence (KL) penalty is used. This yields the overall reward maximised as
@@
where $\beta$ controls the strength of the KL penalty (unless specified otherwise we set $\beta=1\times 10^{-4}$). This procedure, which only collects preferences once in the entire pipeline, has an important disadvantage. Reward models have been found to be poorly calibrated in the higher reward regime bai2022traininghelpfulharmlessassistant and trained policies overoptimise the proxy reward model leading to unstable fine-tuned policies rafailov2024scaling ; gao2022scalinglawsrewardmodel ; ziegler2020finetuninglanguagemodelshuman . Notably, policy optimization induces a divergence between the distributions $\pi_{\theta}(y|x)$ and $\pi^{\text{sft}}(y|x)$. This causes the optimised policy to generate outputs that are different from those seen in the training data $\mathcal{D}$. As a result, the reward model $R_{\phi}$, which was trained on the data $\mathcal{D}$, is now being evaluated on data that it has not seen before (out of distribution).
@@ 3.2 Iterated RLHF and design choices
The problem of the divergence between the distributions $\pi_{\theta}(y|x)$ and $\pi^{sft}(y|x)$ is the one addressed by iterated RLHF. In this process, multiple iterations of steps 1-3 of the RLHF pipeline (namely, collection of preference data, reward model training, and policy optimisation)are repeated as shown in Figure 1. Just as in the single-iteration setting, we start from the checkpoint $\pi^{sft}$ and initialise the reward model from $R^{init}$ with a randomly initialised prediction head. However, there are multiple design choices to be made when choosing how exactly to perform iterated RLHF training. We now describe the process in more detail, highlighting the design choices throughout. Please refer to Algorithm 1 for a schematic of the entire process. For simplicity of notation, we omit explicit references to the policy and reward model parameters $\theta$ and $\phi$, using the subscript $k$ to index iterations instead. During the $k^{th}$ iteration of RLHF, we use the policy from the previous one, denoted by $\pi_{k-1}$ to synthesise pairs of responses for the new preference data denoted by $\tilde{\mathcal{D}}_{k}$.
Indeed, using all policies is unnecessary as it equates to reapplying preference data, but at a higher cost. This new data enables the training of a proxy reward model for which the current policy’s output is in-distribution, potentially mitigating the issue of overoptimisation. Taking into account previous iterations, we now have access to the list of preference data $[\tilde{\mathcal{D}}_{1},...,\tilde{\mathcal{D}}_{k}]$. Here we face the first design choice:
@@
1:Inputs: Prompt dataset $X=\{x_{i}\}_{i=1}^{N}$, $\pi^{sft}$, $R^{init}$, $R^{\star}$, # of iterations $n_{iter}$
2:$\pi_{0}\leftarrow\pi^{sft}$
@@
Combining preference data. Given a list of $k$ preference datasets, the responses in each of which have been generated by different policies $\pi_{1},...,\pi_{k-1}$, we identify three possible options to consolidate them into a single training dataset. The first option (Figure 3.a) is to simply set $\mathcal{D}_{k}=\tilde{\mathcal{D}}_{k}$, only training the reward model on the preference data collected in the current iteration. The second option at the other extreme (i.e., no inter-iteration transfer) is to concatenate all datasets (Figure 3.b). Reusing all the data at each iteration is expected to result in decreased overoptimisation and better approximation with respect to the true reward function. However, this comes with a reward model training computational cost that scales linearly with the number of iterations. Finally, balancing training time and information transfer, we keep the size of the reward model training data constant across iterations by sampling a subsets $\tilde{\mathcal{D}}_{i}$ for $i=1,...,k$ and concatenating the subsets to form $\mathcal{D}_{k}$ (Figure 3.c). Once the training data $\mathcal{D}_{k}$ has been obtained, the proxy reward model $\tilde{R}_{k}$ can be trained on it. $\tilde{R}_{k}$ is initialised from the same base model in all iterations. Having trained the reward model, we now arrive at the second critical design choice:
@@
How do we transfer information from the list of all previously trained proxy reward models $[\tilde{R}_{1},...,\tilde{R}_{k}]$ into a single reward function $R_{k}$ that can be optimised by the policy?
@@
Combining reward models. The reward model is the crucial piece in obtaining generalisable and consistent policies in RLHF, and it is even more important over multiple iterations as effects compound. Given the list $[\tilde{R}_{1},...,\tilde{R}_{k}]$ containing the $k$ proxy reward models leading up to the current iteration the task is to obtain a robust reward function to be optimised. We note that this design choice can be considered in parallel to the combination of preference data, as both target the same outcome of transferring information from previous iterations to the reward function.
@@
To achieve this task we investigate three types of solutions. The first only uses the most recently trained proxy reward model setting $R_{k}=\tilde{R}_{k}$ (Figure 3.d), hence there is no utilisation of previously trained reward models. In contrast, the second option ensembles all previously trained proxy RMs taking the mean of the individual rewards (Figure 3.e) coste2024reward . Since reward model ensembles showed limited improvements in coste2024reward we also evaluate worst-case optimisation (WCO), i.e. optimising the minimum $R_{k}(x,y)=\underset{i=1,...,k}{min}\tilde{R}_{i}(x,y)$. This option comes with the disadvantage of requiring inference on $k$ reward models in parallel. To address the computational cost, we also consider weight averaged reward models (see Figure 3.f) by performing task arithmetic (ilharco2023editingmodelstaskarithmetic, ). More formally, given a sequence of reward models $\tilde{R}_{1},...,\tilde{R}_{k}$, which are parameterised by $\tilde{\phi}_{1},...,\tilde{\phi}_{k}$, respectively, we obtain the proxy reward function $R_{k}$ parameterised by $\phi_{k}$ as follows: The ensemble uses $R_{k}(x,y)=\frac{\sum_{i=1}^{k}\tilde{R}_{i}(x,y)}{k}$ and to obtain the weight averaged reward model we set $\phi_{k}=\frac{\sum_{i=1}^{k}\tilde{\phi}_{i}}{k}$. Having obtained the reward function, the next and final step of each iteration is to optimise it, which leads us to the third and final design choice:
@@
Policy initialisation. The final design choice concerns the initialisation of the policy, i.e., how $\pi^{init}_{k}$ is chosen. The authors of bai2022traininghelpfulharmlessassistant initialise the policy from $\pi^{sft}$ at every iteration, not taking into consideration previously performed computation. We call this initialisation From SFT shown in Figure 3.i. As alternative, we use linear interpolation towards initialisation (LITI) ramé2024warpbenefitsweightaveraged , which was inspired by WiSE-FT proposed by wortsman2022robust . With LITI, shown in Figure3.h, we set $\pi^{init}_{k}=(1-\eta)\pi^{init}_{k-1}+\eta\pi_{k-1}$, where $\eta$ is a hyperparameter balances the optimisation of $R_{k-1}$. Taking $\eta=1$ corresponds to initialising the current policy from the previously fine-tuned one, setting $\pi^{init}_{k}=\pi_{k-1}$. Since continuing fine-tuning of the most recent policy fully relies on the previous iterations, it may suffer from entropy collapse leading to no optimisation in later iterations. Continuing with the fine-tuned policy carries risks if undesirable behaviour learned in previous iterations cannot be unlearned. Note, when performing LITI, the policy is regularised with the KL between the policy and its initialisation $\pi^{init}_{k}$.

### v2

@@ Abstract
Reinforcement learning from human feedback (RLHF) is a widely used method for aligning large language models with human preferences. However, RLHF often suffers from reward model overoptimisation, in which models overfit to the reward function, resulting in non-generalisable policies that exploit the idiosyncrasies and peculiarities of the reward function. A common mitigation is iterated RLHF, in which reward models are repeatedly retrained with updated human feedback and policies are re-optimised. Despite its increasing adoption, the dynamics of overoptimisation in this setting remain poorly understood. In this work, we present the first comprehensive study of overoptimisation in iterated RLHF. We systematically analyse key design choices: how reward model training data is transferred across iterations, which reward function is used for optimisation, and how policies are initialised. Using the controlled AlpacaFarm benchmark, we observe that overoptimisation tends to decrease over successive iterations, as reward models increasingly approximate ground-truth preferences. However, performance gains diminish over time, and while reinitialising from the base policy is robust, it limits optimisation flexibility. Other initialisation strategies often fail to recover from early overoptimisation. These findings offer actionable insights for building more stable and generalisable RLHF pipelines.
@@ 1 Introduction
Reinforcement learning from human feedback (RLHF) has become the standard method for aligning large language models with human preferences (Ziegler et al., 2020; Ouyang et al., 2022; Bai et al., 2022). However, RLHF faces a critical vulnerability: reward model overoptimisation (Gao et al., 2023). As fine-tuning progresses, models learn to overfit to the trained reward function - achieving high scores without genuinely satisfying human intent. This creates brittle policies that exploit loopholes rather than developing robust behaviours, leading to systems that appear aligned during training but fail catastrophically when deployed. Iterated RLHF represents a promising approach to combat this problem. By repeatedly collecting new preferences on the latest policy outputs, retraining the reward model, and fine-tuning the policy (Bai et al., 2022; Xiong et al., 2024), practitioners aim to iteratively close the gap between proxy and true reward. Despite its widespread adoption in industry (Ziegler et al., 2020; Ouyang et al., 2022; Bai et al., 2022), it remains uncertain whether iterated RLHF genuinely resolves overoptimisation, merely postpones the inevitable exploitation of the reward model akin to persistent adversarial policies (Gleave et al., 2020), or perpetuates a recurring cycle of overoptimisation in different forms (Singhal et al., 2024).
@@
Figure 1: The Iterated RLHF framework performs multiple rounds of preference data collection, reward modelling, and policy optimisation. Our research reveals three design choices that dramatically impact performance: (1) how preference data is managed across iterations, (2) which reward function formulation to optimise, and (3) how policies are initialised at each stage. Effectively configuring these elements can significantly reduce overoptimisation.
@@
  • We provide practical guidelines for implementing iterated RLHF, including specific recommendations for preference data management, reward function selection, and policy initialisation strategies.
Figure 2: Iterated RLHF design choices in preference data management, reward function formulation, and policy initialization strongly affect ground truth performance and overoptimisation.
Using a gold-standard reward model to simulate human labellers (Coste et al., 2024; Gao et al., 2023) on the AlpacaFarm dataset (Taori et al., 2023) and working exclusively with open-source models, our experiments yield several key insights: Reward models become increasingly robust across iterations, leading to higher gold reward scores (Figure 2). Performance gains diminish after three iterations for most methods. Concatenating preference data across iterations dramatically outperforms other approaches. Small but persistent overoptimisation remains after four iterations regardless of design choices.
Our results demonstrate that while iterated RLHF significantly improves reward model robustness, it does not fully eliminate overoptimisation. This underscores the need for continued research into more robust alignment methods that can withstand sophisticated specification gaming (Krakovna et al., 2020) by increasingly capable language models.
@@ 2 Related work
RLHF is the standard for aligning large language models to human preference data. The iterated approach has been first discussed by Bai et al. (2022) to fix robustness and calibration issues, attributed to lack of data in the high score regime and has since gained in popularity (Ramé et al., 2024a; Xiong et al., 2024; Ye et al., 2024; Adolphs et al., 2023; Dong et al., 2024; Yuan et al., 2024). Besides training on newly collected preferences, an iterated scheme to train reward models from synthetically generated preference data has been proposed by Wang et al. (2024) and shown to improve performance on the reward model benchmark RewardBench (Lambert et al., 2024), but the authors focus on iterated training of an evaluator and do not study overoptimisation nor the design choices we consider. In the context of Direct Preference Optimisation (DPO) (Rafailov et al., 2023) offline, online and hybrid approaches repeatedly collecting new preference data have been investigated mostly in terms of sample efficiency (Xiong et al., 2024; Das et al., 2024; Muldrew et al., 2024; Mehta et al., 2023). More broadly iterated methods have been investigated for machine teaching (Liu et al., 2017) and to resolve feedback loops caused by model deployment in supervised learning (Perdomo et al., 2020) and also performative RL (Mandal et al., 2023).
Overoptimisation is a common issue in RL, and evidence of this has been frequently reported in the RLHF literature (Ziegler et al., 2020; Stiennon et al., 2020; Gao et al., 2023; Singhal et al., 2024). A promising method to mitigate overoptimisation is using reward model ensembles combined with conservative optimisation (Coste et al., 2024). Several works further explore reward model ensembles in RLHF (Eisenstein et al., 2024; Lou et al., 2024). Notably, Ramé et al. (2024b) introduce weight averaged reward models (WARM) alleviating the inference cost of multiple reward models during training. Following Coste et al. (2024) and Gao et al. (2023) in tackling reward model overoptimisation, several works propose alternative approaches including reward model distillation (Fisch et al., 2024), hidden state regularisation (Yang et al., 2024b), and more (Yang et al., 2024a; Miao et al., 2024; Liu et al., 2024; Gorbatovski et al., 2024). One commonly reported mode of overoptimisation is length bias (Singhal et al., 2024; Park et al., 2024), which can be tackled by disentangling reward signals related to response length from content quality (Chen et al., 2024). To the best of our knowledge, the literature lacks a systematic investigation into overoptimisation in iterated Reinforcement Learning from Human Feedback (RLHF). Such an investigation is not only necessary but also fundamentally important for a deeper understanding and meaningful improvement of fine-tuning methods based on this technique.
@@ 3 Iterated Reinforcement Learning from Human Feedback
In this section, we first outline the process of a single iteration of RLHF and then extend it to the iterated framework. The RLH pipeline consists of the following three steps: 1. Collection of a preference dataset; 2. Reward model training; 3. Policy optimisation on the reward model. Though not an integral part of the RLHF pipeline, it is common in practice for step 1 to be preceded by supervised fine-tuning on labelled examples. To strengthen our investigation we developed a supporting theoretical framework based on performative prediction (Perdomo et al., 2020) that is presented in Appendix A.
@@ 3.1 Single-iteration RLHF
Preference data collection. We start from a supervised fine-tuned policy $\pi^{sft}$ (a policy checkpoint) and use it to collect preference data. The dataset $\mathcal{D}$ contains tuples $\{x_{i},y_{i,0},y_{i,1},p_{i}\}$ for i=1,,Ni=1,...,N, where $x_{i}\in\mathcal{X}$ is a prompt, $y_{i,j}\sim\pi^{sft}(\cdot|x_{i})$ for $j=0,1$ are two responses from $\pi^{sft}$, and $p_{i}$ indicates whether $y_{i,0}$ is preferred over $y_{i,1}$. Following Coste et al. (2024) and Gao et al. (2023), preferences $p_{i}$ are simulated using a gold reward model $R^{\star}$, which is significantly larger in terms of parameter size than the proxy reward models, serving as an approximation for human labels in RLHF.
Reward model training. The proxy reward model $R_{\phi}$ is initialised from model checkpoint $R^{\text{init}}$, with a randomly initialised prediction head, and subsequently trained by minimizing the cross-entropy loss on the preference dataset $\mathcal{D}$. It is standard to use the Bradley-Terry model (Bradley & Terry, 1952), under which the probability of preferring the answer $y_{0}$ over $y_{1}$ given prompt $x$ is given by
@@
Policy optimisation. Having trained the proxy reward model $R_{\phi}$, the policy $\pi_{\theta}$ is initialised from $\pi^{sft}$ and then fine-tuned to optimise $R_{\phi}$. This is commonly achieved with the proximal policy optimization (PPO) algorithm (Schulman et al., 2017). In order to prevent overoptimisation of the proxy reward model and regularise $\pi_{\theta}$ to not diverge too drastically from its initialisation, a Kullback-Leibler divergence (KL) penalty is used. This yields the overall reward maximised as
@@
where $\beta$ controls the strength of the KL penalty (unless specified otherwise we set $\beta=1\times 10^{-4}$). This procedure, which only collects preferences once in the entire pipeline, has an important disadvantage. Reward models have been found to be poorly calibrated in the higher reward regime (Bai et al., 2022) and trained policies overoptimise the proxy reward model leading to unstable fine-tuned policies (Rafailov et al., 2024; Gao et al., 2023; Ziegler et al., 2020). Notably, policy optimization induces a divergence between the distributions $\pi_{\theta}(y|x)$ and $\pi^{\text{sft}}(y|x)$. This causes the optimised policy to generate outputs that are different from those seen in the training data $\mathcal{D}$. As a result, the reward model $R_{\phi}$, which was trained on the data $\mathcal{D}$, is now being evaluated on data that it has not seen before (out of distribution).
@@ 3.2 Iterated RLHF and design choices
The problem of the divergence between the distributions $\pi_{\theta}(y|x)$ and $\pi^{sft}(y|x)$ is the one addressed by iterated RLHF. In this process, multiple iterations of steps 1-3 of the RLHF pipeline (namely, collection of preference data, reward model training, and policy optimisation)are repeated as shown in Figure 1. Just as in the single-iteration setting, we start from the checkpoint $\pi^{sft}$ and initialise the reward model from $R^{init}$ with a randomly initialised prediction head. However, there are multiple design choices to be made when choosing how exactly to perform iterated RLHF training. We now describe the process in more detail, highlighting the design choices throughout. Please refer to Algorithm˜1 for a schematic of the entire process. For simplicity of notation, we omit explicit references to the policy and reward model parameters $\theta$ and $\phi$, using the subscript $k$ to index iterations instead. During the $k^{th}$ iteration of RLHF, we use the policy from the previous one, denoted by $\pi_{k-1}$ to synthesise pairs of responses for the new preference data denoted by $\tilde{\mathcal{D}}_{k}$.
Indeed, using all policies is unnecessary as it equates to reapplying preference data, but at a higher cost. This new data enables the training of a proxy reward model for which the current policy’s output is in-distribution, potentially mitigating the issue of overoptimisation. Taking into account previous iterations, we have access to the list of preference data $\tilde{\mathcal{D}}_{1},...,\tilde{\mathcal{D}}_{k}]$. Here we face the first design choice:
@@
1:Inputs: Prompt dataset $X=\{x_{i}\}_{i=1}^{N}$, $\pi^{sft}$, $^{init}, RR^{\star}$, # of iterations $n_{iter}$
[…]
@@
Combining preference data. Given a list of $k$ preference datasets, the responses in each of which have been generated by different policies $\pi_{1},...,\pi_{k-1}$, we identify three possible options to consolidate them into a single training dataset. The first option (Figure 3.a) is to simply set $\mathcal{D}_{k}=\tilde{\mathcal{D}}_{k}$, only training the reward model on the preference data collected in the current iteration. The second option at the other extreme (i.e., no inter-iteration transfer) is to concatenate all datasets (Figure 3.b). Reusing all the data at each iteration is expected to result in decreased overoptimisation and better approximation with respect to the true reward function. However, this comes with a reward model training computational cost that scales linearly with the number of iterations. Finally, balancing training time and information transfer, we keep the size of the reward model training data constant across iterations by sampling a subsets $\tilde{\mathcal{D}}_{i}$ for i=1,,ki=1,...,k and concatenating the subsets to form $\mathcal{D}_{k}$ (Figure 3.c). Once the training data $\mathcal{D}_{k}$ has been obtained, the proxy reward model $\tilde{R}_{k}$ can be trained on it. $\tilde{R}_{k}$ is initialised from the same base model in all iterations. Having trained the reward model, we now arrive at the second critical design choice:
@@
How do we transfer information from the list of all previously trained proxy reward models $\tilde{R}_{1},...,\tilde{R}_{k}]$ into a single reward function $R_{k}$ that can be optimised by the policy?
@@
Combining reward models. The reward model is the crucial piece in obtaining generalisable and consistent policies in RLHF, and it is even more important over multiple iterations as effects compound. Given the list $\tilde{R}_{1},...,\tilde{R}_{k}]$ containing the $k$ proxy reward models leading up to the current iteration the task is to obtain a robust reward function to be optimised. We note that this design choice can be considered in parallel to the combination of preference data, as both target the same outcome of transferring information from previous iterations to the reward function.
@@
To achieve this task we investigate three types of solutions. The first only uses the most recently trained proxy reward model setting $R_{k}=\tilde{R}_{k}$ (Figure 3.d), hence there is no utilisation of previously trained reward models. In contrast, the second option ensembles all previously trained proxy RMs taking the mean of the individual rewards (Figure 3.e) (Coste et al., 2024). Since reward model ensembles showed limited improvements in Coste et al. (2024) we also evaluate worst-case optimisation (WCO), i.e., optimising the minimum $R_{k}(x,y)=\underset{i=1,...,k}{min}\tilde{R}_{i}(x,y)$. This option comes with the disadvantage of requiring inference on $k$ reward models in parallel. To address the computational cost, we also consider weight averaged reward models (see Figure 3.f) by performing task arithmetic (Ilharco et al., 2023). More formally, given a sequence of reward models $\tilde{R}_{1},...,\tilde{R}_{k}$, which are parameterised by $\tilde{\phi}_{1},...,\tilde{\phi}_{k}$, respectively, we obtain the proxy reward function $R_{k}$ parameterised by $\phi_{k}$ as follows: The ensemble uses $R_{k}(x,y)=\frac{\sum_{i=1}^{k}\tilde{R}_{i}(x,y)}{k}$ and to obtain the weight averaged reward model we set $\phi_{k}=\frac{\sum_{i=1}^{k}\tilde{\phi}_{i}}{k}$. Having obtained the reward function, the next and final step of each iteration is to optimise it, which leads us to the third and final design choice:
@@
Policy initialisation. The final design choice concerns the initialisation of the policy, i.e., how $\pi^{init}_{k}$ is chosen. Bai et al. (2022) initialise the policy from $\pi^{sft}$ at every iteration, not taking into consideration previously performed computation. We call this initialisation From SFT shown in Figure 3.i. As alternative, we use linear interpolation towards initialisation (LITI) (Ramé et al., 2024a), which was inspired by WiSE-FT proposed by (Wortsman et al., 2022). With LITI, shown in Figure3.h, we set $\pi^{init}_{k}=(1-\eta)\pi^{init}_{k-1}+\eta\pi_{k-1}$, where $\eta$ is a hyperparameter balances the optimisation of $R_{k-1}$. Taking $\eta=1$ corresponds to initialising the current policy from the previously fine-tuned one, setting $\pi^{init}_{k}=\pi_{k-1}$. Since continuing fine-tuning of the most recent policy fully relies on the previous iterations, it may suffer from entropy collapse leading to no optimisation in later iterations. Continuing with the fine-tuned policy carries risks if undesirable behaviour learned in previous iterations cannot be unlearned. Note, when performing LITI, the policy is regularised with the KL between the policy and its initialisation $\pi^{init}_{k}$.